# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

### v1

@@ 3.3. Evaluation & Config Modules
The evaluation module provides an evaluation() interface to measure the performance of MapRL models on downstream tasks. the VecCity library offers a range of metrics tailored to different tasks. Specifically, it includes MAE, MSE, RMSE, R2, and MAPE for regression tasks; Accuracy, Precision@k, Recall@k, F1@k, and Mean Rank (MR) for classification tasks; and Accuracy@k and MR for trajectory similarity search tasks.
The config module provides a configuration file as an interface for users to arrange the “pre-training $\rightarrow$ fine-tuning $\rightarrow$ evaluation” pipeline. To train and test a model using VecCity, users simply specify the required pre-built or custom models and functions for each step in the pipeline within the configuration file. The library then automatically generates the necessary script to run experiments based on this configuration. Users only need to execute the script, and VecCity will invoke the interface functions of each module to complete the experiment.
@@ 4. Performance Benchmark
In this section, we present performance comparison experiments for the pre-built MapRL models in VecCity. These experiments serve three purposes: $i$) Demonstrating Usability and Power: To show that VecCity is a user-friendly and powerful toolkit for MapRL model development and evaluation. $ii$) Establishing Performance Benchmarks: To provide standard performance benchmarks for mainstream MapRL models using the datasets pre-loaded in VecCity. $iii$) Analyzing Component Contributions: To explore the contributions of different components in existing MapRL models, offering insights for designing more effective and robust models.
@@ 4.1.1. Datasets
The dataset selection for our experiments follows these guiding principles: for each type of MapRL model, we incorporate all datasets that meet the task requirements. Specifically, for POI-oriented MapRL models, we use datasets from New York City, Chicago, Tokyo, Singapore, and San Francisco. For road-segment-oriented models, the datasets include Beijing, Xi’an, Chengdu, Porto, and San Francisco. For land-parcel-oriented models, we adopt datasets from Beijing, Xi’an, Chengdu, Porto, and San Francisco. Each dataset is divided into training, validation, and test sets following a 6:2:2 ratio, ensuring balanced and reliable performance evaluation. For each downstream task, we apply end-to-end fine-tuning to achieve optimal performance.
@@ 4.1.2. Experimental Settings
In our experiments, the dimension of map entity representation vectors across all models is set to 128. For other model configurations, such as the number of layers and the dimensions of intermediate variables, we adopt the default settings from the original papers as initial values. We then use a trial-and-error approach to tune the parameters on each dataset, searching for the best performance configurations. All experiments are executed on an Ubuntu 20.04 system equipped with NVIDIA GeForce 3090 GPUs. The default batch size during model training is set to 64. In cases of out-of-memory (OOM) issues, the batch size is halved iteratively, with a minimum batch size of 8.
@@
Dataset Beijing Chengdu Xi’an Porto San Francisco Overall
Task ASI TTE ASI TTE ASI TTE ASI TTE ASI TTE Per Ave Rank Per Rank
Metric MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓
RN2Vec 3.566 5.744 656.6 3050 6.490 13.90 105.9 154.8 6.237 12.72 130.38 190.5 4.380 7.896 97.8 144.5 2.978 5.797 289.0 1479 6.0 7
HRNR 3.157 5.508 625.1 3066 6.364 13.66 100.26 149.78 5.672 11.11 129.24 191.1 4.357 7.867 104.5 150.2 2.510 5.374 287.8 1445 5.8 6
SARN 2.798 5.479 566.8 2578 6.025 13.41 95.64 148.3 5.315 10.60 117.0 178.7 4.286 7.779 99.5 145.8 2.481 5.303 279.0 1502 4.1 5
Toast 2.837 5.512 513.0 2499 6.377 13.62 90.7 145.4 5.619 10.96 116.3 184.9 4.354 7.883 96.4 144.6 2.553 5.295 246.4 1270 3.6 4
HRoad 2.726 5.411 610.3 2527 5.970 13.22 96.24 146.6 5.537 10.81 116.6 187.4 4.212 7.693 97.2 144.3 2.476 5.250 287.6 1546 3.5 3
START 2.738 5.445 509.4 2489 6.066 13.47 92.3 145.9 5.503 10.92 116.4 183.4 4.210 7.792 95.8 145.7 2.450 5.333 261.8 1431 2.6 2
JCLRNT 2.771 5.466 475.9 2442 6.354 13.27 90.2 142.7 5.660 11.02 115.9 183.5 4.267 7.699 94.6 143.4 2.534 5.368 264.5 1416 2.4 1
Task STS Efficiency STS Efficiency STS Efficiency STS Efficiency STS Efficiency Eff Ave Rank Eff Rank
Metric ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓
RN2Vec 0.789 10.53 5.192 331 0.631 45.32 0.708 19 0.695 11.208 0.826 16 0.893 8.296 1.453 3 0.976 6.179 3.524 35 2.7 2
HRNR 0.809 9.915 5.699 281 0.698 37.12 1.188 5 0.709 10.580 1.321 4 0.807 11.170 1.948 80 0.915 6.988 4.066 23 3.6 4
SARN 0.892 9.176 0.570 30 0.563 40.31 0.549 5 0.686 7.384 0.551 5 0.883 8.753 0.551 10 0.963 6.879 0.563 19 1.2 1
Toast 0.828 9.755 10.01 495 0.678 18.64 1.840 192 0.749 5.874 1.566 132 0.895 9.031 3.045 252 0.978 5.934 7.021 253 6.0 6
HRoad 0.947 8.491 5.398 66 0.583 32.05 0.913 6 0.736 7.122 1.031 5 0.895 10.152 1.660 11 0.947 8.491 3.728 44 3.3 3
START 0.948 7.867 6.238 1719 0.713 15.59 2.139 582 0.755 5.411 2.002 312 0.905 6.702 2.745 735 0.934 6.828 4.743 783 6.7 7
JCLRNT 0.879 9.034 5.114 434 0.724 14.34 1.049 287 0.797 4.746 0.912 150 0.927 4.778 1.649 345 0.957 6.284 3.630 198 4.3 5
4.2. Overall Experimental Results
@@ 4.2.1. Experimental Results for POI MapRL
Table 3 presents the performance of MapRL models on downstream tasks, including POI Classification (POIC), Next POI Prediction (NPP), and Trajectory User Link (TUL). Each task is evaluated using two types of metrics. Additionally, model efficiency is assessed with two indicators: the number of parameters (Param) in millions and the time cost per epoch during pre-training (Speed) in seconds. The table also reports average performance rankings across tasks (Pre Ave Rank) and the ranking based on the models’ average performance (Pre Rank). Similarly, efficiency rankings are given by Eff Ave Rank and Eff Rank. From Tab. 3, we have the following observations.
@@
1) Importance of Sequence-based Encoders: Models combining token-based and sequence-based encoders – such as CTLE, CACSR, and Hier – consistently outperform those relying solely on token-based encoders, highlighting the value of trajectory auxiliary data. Among these, models using Augmented Trajectory Contrastive Learning (ATrCL) and Masked Trajectory Recovery (MTR) tasks (e.g., , CTLE and CACSR) achieve better performance than those focused on Trajectory Prediction (TrajP) tasks (e.g., , Hier). This is likely because ATrCL and MTR capture long-range temporal dependencies across entire trajectories, while TrajP focuses only on consecutive POI dependencies.
@@
2) Impact of Relation Networks: For token-based models, those incorporating relation networks – such as Teaser and P2Vec – outperform models without them, such as SkipGram and Tale, underscoring the value of relation network auxiliary data. However, for models with both token-based and sequence-based encoders, the inclusion of relation networks offers little additional benefit. This is likely due to the sparsity of POI trajectories, where the number of trajectories between two POIs is often dictated by geographical proximity. As a result, the information provided by relation networks overlaps significantly with that from trajectory data. Once the dependency within the trajectory data has been fully exploited, the relational network contributes limited additional information.
@@
3) TokRI vs. TRCL Tasks: Token-based models using the Token Relation Inference (TokRI) task, such as Tale and P2Vec, relatively outperform those employing the Token Relation Contrastive Learning (TRCL) task, such as SkipGram and Teaser. With sufficient data, supervised learning via TokRI effectively leverages labeled data to generate more discriminative feature representations. In contrast, the indirect nature of contrastive learning in TRCL offers less direct utilization of labeled information, resulting in comparatively lower model performance.
4) Efficiency of Token-based Models: Token-based models exhibit lower computational complexity, making them more efficient. This makes them especially suitable for application scenarios where computational resources are limited, and performance requirements can be moderately compromised.

### v2

@@ 3.3. Evaluation & Config Modules
The evaluation module provides an evaluation() interface to assess MapRL model performance across downstream tasks. the VecCity library supports various metrics tailored to different task types: MAE, MSE, RMSE, R2, and MAPE for regression tasks; Accuracy@k, Recall@k, F1@k, and Mean Rank (MR) for classification tasks; and Accuracy@k and MR for trajectory similarity search.
The config module serves as the control center, allowing users to configure the entire “pre-training $\rightarrow$ fine-tuning $\rightarrow$ evaluation” pipeline through a configuration file. To train and evaluate a model with VecCity, users simply specify the required pre-built or custom models and functions in the configuration file. The library then automatically generates the necessary script to run experiments. Users only need to execute the script, and VecCity will invoke the interface functions of each module to complete the experiment.
@@ 4. Performance Benchmark
In this section, we present performance comparison experiments for the pre-built MapRL models in VecCity. These experiments serve three purposes: $i$) Demonstrating Usability and Power: To show that VecCity is a user-friendly and powerful toolkit for MapRL model development and evaluation. $ii$) Establishing Performance Benchmarks: To provide standard performance comparisons of mainstream MapRL models using the datasets pre-loaded in VecCity. $iii$) Analyzing Model Efficiency. Evaluating the computational efficiency of different MapRL models to offer insights for designing more efficient approaches.
@@ 4.1.1. Datasets
The dataset selection for our experiments follows these guiding principles: for each type of MapRL model, we incorporate all datasets that meet the task requirements. Specifically, for POI-oriented MapRL models, we use datasets from New York, Chicago, Tokyo, Singapore, and San Francisco. For road-segment-oriented models, the datasets include Beijing, Chengdu, Xi’an, Porto, and San Francisco. For land-parcel-oriented models, we use datasets from Beijing, Chengdu, Xi’an, Porto, and San Francisco. Each dataset is divided into training, validation, and test sets following a 6:2:2 ratio, ensuring a balanced and reliable performance evaluation. For each downstream task, we apply end-to-end fine-tuning to achieve optimal performance.
@@ 4.1.2. Experimental Settings
In our experiments, the dimension of map entity representation vectors across all models is set to 128. For other model configurations, such as the number of layers and the dimensions of intermediate variables, we adopt the default settings from the original papers as initial values. We then use a trial-and-error approach to tune the parameters on each dataset, searching for the best performance configurations. Each model is run five times with different seeds, and the results are reported as the mean and standard deviation. All experiments are executed on an Ubuntu 20.04 system equipped with NVIDIA GeForce 3090 GPUs. The default batch size during model training is set to 64. In cases of out-of-memory (OOM) issues, the batch size is halved iteratively, with a minimum batch size of 8.
Table 3. Performance comparison of POI MapRL methods across various datasets (Mean ± Std). Bold and underlined text indicates the best and second-best results, respectively. ’↑’ denotes that a higher value is better, while ’↓’ indicates that a lower value is preferred.
Data Task Metric SkipG Tale Teaser Hier P2Vec CACSR CTLE
New York POIC ACC@1↑ 0.07±0.00 0.06±0.01 0.07±0.01 0.07±0.01 0.08±0.01 0.06±0.02 0.06±0.01
F1↑ 0.01±0.00 0.02±0.01 0.01±0.01 0.01±0.00 0.02±0.01 0.01±0.00 0.01±0.00
NPP ACC@1↑ 0.10±0.02 0.14±0.01 0.10±0.02 0.12±0.00 0.13±0.01 0.14±0.05 0.15±0.01
ACC@5↑ 0.25±0.04 0.33±0.00 0.25±0.05 0.29±0.01 0.16±0.00 0.33±0.04 0.33±0.01
TUL ACC@1↑ 0.57±0.03 0.67±0.00 0.57±0.02 0.58±0.03 0.64±0.01 0.64±0.03 0.69±0.00
F1↑ 0.26±0.03 0.44±0.01 0.27±0.03 0.28±0.04 0.33±0.00 0.39±0.07 0.45±0.01
Tokyo POIC ACC@1↑ 0.35±0.01 0.22±0.00 0.38±0.02 0.34±0.04 0.32±0.00 0.30±0.04 0.30±0.04
F1↑ 0.02±0.01 0.05±0.01 0.02±0.01 0.01±0.01 0.04±0.00 0.02±0.00 0.02±0.00
NPP ACC@1↑ 0.16±0.00 0.16±0.01 0.16±0.01 0.16±0.00 0.16±0.01 0.16±0.06 0.16±0.00
ACC@5↑ 0.32±0.01 0.30±0.00 0.30±0.04 0.34±0.00 0.33±0.01 0.33±0.01 0.33±0.01
TUL ACC@1↑ 0.43±0.01 0.40±0.01 0.44±0.02 0.44±0.03 0.45±0.00 0.46±0.10 0.50±0.00
F1↑ 0.23±0.01 0.23±0.01 0.23±0.01 0.23±0.01 0.25±0.00 0.28±0.10 0.29±0.00
Chicago POIC ACC@1↑ 0.07±0.00 0.10±0.00 0.07±0.01 0.08±0.00 0.12±0.01 0.10±0.01 0.10±0.02
F1↑ 0.01±0.00 0.02±0.01 0.00±0.00 0.01±0.00 0.02±0.01 0.02±0.01 0.02±0.01
NPP ACC@1↑ 0.08±0.02 0.10±0.01 0.06±0.03 0.10±0.03 0.15±0.01 0.15±0.05 0.18±0.01
ACC@5↑ 0.24±0.05 0.24±0.01 0.18±0.08 0.26±0.06 0.36±0.01 0.40±0.04 0.38±0.01
TUL ACC@1↑ 0.46±0.08 0.62±0.01 0.43±0.07 0.55±0.02 0.66±0.01 0.67±0.07 0.76±0.01
F1↑ 0.20±0.07 0.37±0.01 0.19±0.01 0.30±0.01 0.42±0.01 0.44±0.06 0.60±0.03
Singapore POIC ACC@1↑ 0.09±0.04 0.04±0.01 0.06±0.00 0.12±0.04 0.07±0.00 0.10±0.04 0.13±0.03
F1↑ 0.00±0.00 0.01±0.01 0.01±0.01 0.00±0.00 0.03±0.01 0.01±0.00 0.00±0.00
NPP ACC@1↑ 0.03±0.00 0.05±0.01 0.03±0.00 0.05±0.02 0.08±0.01 0.11±0.04 0.10±0.03
ACC@5↑ 0.06±0.01 0.10±0.00 0.07±0.01 0.11±0.03 0.18±0.01 0.18±0.05 0.14±0.05
TUL ACC@1↑ 0.26±0.06 0.31±0.01 0.27±0.06 0.27±0.08 0.40±0.01 0.40±0.05 0.44±0.03
F1↑ 0.11±0.04 0.17±0.01 0.14±0.01 0.16±0.01 0.18±0.01 0.18±0.07 0.25±0.00
San Francisco POIC ACC@1↑ 0.06±0.03 0.01±0.00 0.06±0.01 0.07±0.02 0.04±0.01 0.06±0.02 0.06±0.01
F1↑ 0.00±0.00 0.01±0.01 0.00±0.00 0.00±0.00 0.02±0.01 0.01±0.01 0.00±0.00
NPP ACC@1↑ 0.02±0.01 0.01±0.00 0.02±0.00 0.04±0.01 0.03±0.01 0.03±0.03 0.06±0.00
ACC@5↑ 0.06±0.01 0.06±0.01 0.06±0.01 0.10±0.01 0.06±0.01 0.06±0.06 0.12±0.01
TUL ACC@1↑ 0.25±0.08 0.34±0.01 0.30±0.11 0.35±0.12 0.52±0.01 0.53±0.06 0.61±0.01
F1↑ 0.13±0.01 0.15±0.00 0.10±0.06 0.15±0.05 0.26±0.01 0.29±0.02 0.32±0.01
Per. Avg. Rank 5.8 5.7 4.5 4.2 3.0 2.9 2.4
4.2. Overall Experimental Results
@@
Data Task Metric RN2Vec HRNR SARN Toast HRoad START JCLRNT
Beijing ASI MAE↓ 3.58±0.01 3.16±0.00 2.81±0.00 2.74±0.01 2.71±0.01 2.72±0.01 2.80±0.02
RMSE↓ 5.73±0.00 5.59±0.00 5.48±0.00 5.51±0.01 5.41±0.00 5.39±0.00 5.48±0.01
TTE MAE↓ 661±14.0 631±19.3 596±20.8 533±33.0 585±16.3 512±19.9 514±38.3
RMSE↓ 2978±48 2968±42 2581±84 2576±79 2554±35 2502±37 2464±36
STS ACC@3↑ 0.76±0.03 0.82±0.03 0.84±0.03 0.79±0.10 0.87±0.01 0.90±0.03 0.90±0.04
MR↓ 12.4±1.54 11.5±1.94 13.30±2.17 12.8±2.81 14.0±1.42 10.9±1.98 10.4±1.15
Chendu ASI MAE↓ 6.57±0.06 6.42±0.04 5.99±0.02 6.36±0.03 5.93±0.02 6.05±0.01 6.35±0.08
RMSE↓ 14.0±0.07 13.7±0.07 13.3±0.04 13.3±0.16 13.2±0.00 13.5±0.01 13.5±0.13
TTE MAE↓ 106±0.06 100±0.04 88.3±4.27 95.0±2.44 88.5±5.48 95.8±3.12 97.0±4.08
RMSE↓ 155±0.07 150±0.05 157±6.70 150±2.75 137±5.67 151±3.84 148±3.79
STS ACC@3↑ 0.69±0.04 0.73±0.06 0.68±0.13 0.69±0.02 0.68±0.05 0.76±0.03 0.72±0.00
MR↓ 48.1±1.67 43.0±3.37 36.6±3.14 22.4±3.66 24.4±3.01 19.2±2.72 18.7±2.53
Xi’an ASI MAE↓ 6.33±0.03 5.79±0.12 5.36±0.03 5.63±0.01 5.41±0.07 5.38±0.07 5.67±0.02
RMSE↓ 13.1±0.37 11.2±0.10 10.6±0.01 10.8±0.12 10.7±0.05 10.8±0.07 10.9±0.06
TTE MAE↓ 137±6.67 134±5.09 110±4.26 122±4.52 110±3.68 115±15.4 108±6.82
RMSE↓ 199±8.23 203±11.9 169±5.63 191±3.62 172±8.94 170±13.9 178±8.34
STS ACC@3↑ 0.71±0.01 0.74±0.04 0.72±0.07 0.80±0.03 0.73±0.01 0.75±0.04 0.77±0.02
MR↓ 12.9±1.71 12.1±1.48 12.5±3.62 16.5±4.28 19.0±2.95 10.8±3.54 11.9±3.15
Porto ASI MAE↓ 4.39±0.00 4.40±0.00 4.29±0.00 4.36±0.01 4.20±0.01 4.21±0.00 4.35±0.05
RMSE↓ 7.90±0.00 7.58±0.00 7.78±0.00 7.93±0.05 7.71±0.01 7.79±0.04 7.79±0.05
TTE MAE↓ 99.5±1.78 101±1.98 97.2±1.38 98.4±1.81 98.5±2.83 97.5±6.92 96.8±1.30
RMSE↓ 150±3.60 147±2.60 143±1.80 147±2.48 140±2.22 158±7.04 145±1.03
STS ACC@3↑ 0.80±0.10 0.82±0.11 0.83±0.03 0.90±0.06 0.83±0.04 0.88±0.02 0.87±0.04
MR↓ 14.3±1.91 12.0±1.81 17.9±3.40 12.1±4.03 17.2±1.80 13.7±2.44 11.4+3.38
San Francisco ASI MAE↓ 2.98±0.00 2.51±0.00 2.49±0.01 2.56±0.00 2.48±0.01 2.45±0.00 2.42±0.01
RMSE↓ 5.80±0.00 5.38±0.01 5.29±0.01 5.29±0.00 5.25±0.00 5.33±0.00 5.37±0.00
TTE MAE↓ 289±1.02 288±1.01 279±3.21 274±9.01 270±7.61 268±10.2 275±9.30
RMSE↓ 1479±4.3 1449±4.4 1499±3.6 1372±28 1451±16 1497±38 1446±21
STS ACC@3↑ 0.84±0.04 0.91±0.00 0.86±0.07 0.94±0.04 0.92±0.01 0.90±0.07 0.87±0.05
MR↓ 7.20±3.19 7.00±0.01 8.48±2.05 9.22±3.29 6.55±1.22 6.71±2.01 6.49±1.63
Per Avg Rank 6.3 4.9 3.9 3.9 3.2 2.9 2.8
@@ 4.2.1. Experimental Results for POI MapRL
Table 3 shows the performance of MapRL models on three downstream tasks: POI Classification (POIC), Next POI Prediction (NPP), and Trajectory User Link (TUL), each evaluated using two types of metrics. The table also reports average performance rankings across tasks (Pre Ave Rank) and the ranking based on the models’ average performance (Pre Rank). From Tab. 3, we have the following observations.
@@
1) Importance of Sequence-based Encoders: Models integrating token- and sequence-based encoders (e.g., , CTLE, CACSR, Hier) outperform token-based models, indicating the value of trajectory data. Among these, models using Augmented Trajectory Contrastive Learning (ATrCL) and Masked Trajectory Recovery (MTR) tasks (e.g., , CTLE, CACSR) achieve better performance than those focused on Trajectory Prediction (TrajP) tasks (e.g., , Hier). This is likely because ATrCL and MTR capture long-range temporal dependencies across entire trajectories, while TrajP focuses only on consecutive POI dependencies.
@@
2) Impact of Relation Networks: For token-based models, those incorporating relation networks (e.g., , Teaser, P2Vec) outperform models without them (e.g., , SkipGram, Tale), highlighting their value. However, for models combining token- and sequence-based encoders, relation networks provide little benefit. This is likely due to the sparsity of POI trajectories, where the number of trajectories between two POIs is often dictated by geographical proximity. As a result, the information provided by relation networks overlaps significantly with that from trajectory data. Once the dependency within the trajectory data has been fully exploited, the relational network contributes limited additional information.
@@
3) TokRI vs. TRCL Tasks: Token-based models utilizing the Token Relation Inference (TokRI) task, such as P2Vec, generally outperform those employing Token Relation Contrastive Learning (TRCL), such as SkipGram and Teaser. With sufficient data, supervised learning via TokRI effectively leverages labeled data to generate more discriminative representations. In contrast, the indirect nature of contrastive learning in TRCL offers less direct utilization of labeled information, resulting in comparatively lower model performance.
4) Instability Introduced by Contrastive Learning. Models that rely on contrastive learning tasks, such as SkipGram, Teaser, and CACSR, exhibit instability due to their sensitivity to dataset partitioning and negative sample selection. In contrast, P2Vec and Tale, which use TokRI—a supervised pretraining task—demonstrate greater stability by explicitly modeling relationships between map entities.
Data Task Metric ZEMob GMEL HDGE MVURE MGFN ReMVC HREP
Beijing LPC ACC@1↑ 0.43±0.03 0.46±0.03 0.41±0.00 0.48±0.00 0.59±0.02 0.75±0.02 0.50±0.01
F1↑ 0.11±0.00 0.61±0.03 0.22±0.01 0.28±0.01 0.65±0.04 0.71±0.02 0.38±0.01
FI MAE↓ 30.9±0.03 27.7±0.01 20.8±2.91 29.6±0.83 30.1±0.02 27.8±0.02 20.6±0.33
RMSE↓ 55.3±0.04 59.0±0.04 44.8±2.79 53.7±1.36 54.5±0.02 52.7±0.04 42.6±0.86
MI MAE↓ 0.43±0.03 0.46±0.03 0.34±0.03 0.27±0.01 0.30±0.03 0.30±0.01 0.32±0.00
RMSE↓ 0.80±0.02 0.74±0.02 0.70±0.03 0.71±0.02 0.76±0.04 0.69±0.02 0.66±0.01
Chengdu LPC ACC@1↑ 0.36±0.01 0.38±0.04 0.42±0.03 0.48±0.00 0.48±0.08 0.66±0.03 0.54±0.01
F1↑ 0.17±0.03 0.16±0.03 0.27±0.06 0.27±0.01 0.35±0.12 0.59±0.03 0.46±0.01
FI MAE↓ 155±0.03 155±0.02 159±2.02 156±3.10 158±5.69 148±0.02 156±2.8
RMSE↓ 441±0.02 441±0.04 403±1.12 443±4.28 337±24.4 431±0.03 396±9.7
MI MAE↓ 3.64±0.03 4.62±0.04 5.17±0.01 3.97±0.71 5.14±0.02 5.29±0.04 5.47±0.12
RMSE↓ 29.7±0.02 29.7±0.03 28.7±0.15 29.6±0.12 29.3±1.14 29.1±0.06 28.3±0.07
Xi’an LPC ACC@1↑ 0.36±0.03 0.36±0.03 0.44±0.02 0.53±0.01 0.48±0.09 0.57±0.00 0.59±0.01
F1↑ 0.19±0.04 0.42±0.03 0.25±0.01 0.26±0.02 0.28±0.07 0.46±0.03 0.45±0.02
FI MAE↓ 134±0.03 133±0.02 124±5.87 127±3.85 119±9.30 132±0.04 129±0.80
RMSE↓ 326±0.02 324±0.03 289±26.3 318±4.77 241±28.4 325±0.03 297±9.10
MI MAE↓ 3.83±0.03 3.71±0.04 3.92±0.03 3.33±0.35 3.70±0.11 3.91±0.05 4.02±0.04
RMSE↓ 17.7±0.03 17.6±0.03 16.6±0.14 17.4±0.00 16.8±1.32 16.8±0.05 16.7±0.20
Porto LPC ACC@1↑ 0.34±0.04 0.33±0.04 0.42±0.05 0.30±0.01 0.30±0.01 0.44±0.03 0.36±0.02
F1↑ 0.21±0.01 0.38±0.03 0.41±0.07 0.27±0.02 0.20±0.03 0.41±0.04 0.32±0.03
FI MAE↓ 528±0.03 525±0.01 501±18.0 402±21.8 477±37.3 527±0.01 289±12.1
RMSE↓ 595±0.02 592±0.04 582±14.4 472±21.1 443±36.8 595±0.03 355±27.3
MI MAE↓ 44.9±0.02 41.1±0.03 30.8±4.57 37.8±0.65 29.8±7.26 39.4±0.05 28.3±5.30
RMSE↓ 72.3±0.02 56.2±0.04 45.7±4.91 54.1±4.68 45.5±9.79 56.7±0.03 42.2±8.93
San Francisco LPC ACC@1↑ 0.84±0.02 0.64±0.01 0.82±0.01 0.84±0.02 0.83±0.01 0.86±0.03 0.84±0.01
F1↑ 0.53±0.03 0.56±0.01 0.50±0.03 0.52±0.05 0.44±0.10 0.47±0.04 0.57±0.03
FI MAE↓ 509±0.04 514±0.01 387±23.1 349±23.7 356±46 445±0.01 324±25.8
RMSE↓ 910±0.04 917±0.02 693±35.8 700±44.2 748±80 852±0.02 624±31
MI MAE↓ 5.88±0.03 7.40±0.02 6.51±0.51 6.46±0.42 6.40±0.84 6.87±0.04 5.89±0.50
RMSE↓ 23.5±0.01 22.3±0.01 21.5±0.28 20.6±1.21 20.2±2.34 22.7±0.03 19.1±1.51
Per Avg Rank 5.67 5.10 4.00 3.73 3.53 3.50 2.43